# Imaging a semi-classical horizonless compact object with strong redshift

### v1

@@ V Physical QHCO images
Although it is impossible to distinguish these two sets of images by the naked eye, some subtle differences will appear when there are light emissions near the ``would-be horizon" at $r=a_{0}$, e.g. the GLM1 and GLM3 emission models, since the photons emitted there may be observable. Because the direct emission of the event horizon (or the would-be horizon for QHCOs) corresponds to the inner shadow in the images, we can expect that some subtle differences between the darkened QHCO images and the Schwarzschild images would appear near their inner shadows, and they would originate from the photons emitted in the near-surface region with width $\Delta r\sim\sigma/a_{0}$.

### v2

@@ V Physical QHCO images
Although it is impossible to distinguish these two sets of images by the naked eye, some subtle differences will appear when there are light emissions near the “would-be horizon” at $r=a_{0}$, e.g. the GLM1 and GLM3 emission models, since the photons emitted there may be observable. Because the direct emission of the event horizon (or the would-be horizon for QHCOs) corresponds to the inner shadow in the images, we can expect that some subtle differences between the darkened QHCO images and the Schwarzschild images would appear near their inner shadows, and they would originate from the photons emitted in the near-surface region with width $\Delta r\sim\sigma/a_{0}$.